[clas_members] [New Logentry] RC Daily summary 6 Sept

weinstei at jlab.org weinstei at jlab.org
Tue Sep 6 13:50:01 EDT 2022

Logentry Text:
We changed to the carbon target at 1630 yesterday and then had difficulty bringing beam to the hall due to bleed thru.  When the Hall B laser was off and the raster was on, it would trip the halo counters.  We determined the problem was from Hall C.

Starting around 21:00, Eugene, the RC, and the shift expert started with a small raster pattern and slowly increased it until we had decent beam.  We realized overnight that the 1H00 steering coils were energized.  At 0900 this morning, Eugene, the RC, and the shift expert had MCC turn off the 1h00 steering coils.  We then re-centered the beam and expanded the raster pattern.

1200 access for target change to ND3 target.

** RC meeting at 1400 today **

Sorry for the late notice.

- Larry Weinstein


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4038431
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220906/12f3705e/attachment.html>

More information about the clas_members mailing list